OpenStudy (anonymous):

Solve by elimination method. 5x+3y=-7 7x-2y=19 I know that y= 43/31, but I do not know what I am supposed to do when I plg 43/31 in for x in the equation... the answer they came up with is ; 43/31, -144/31 How do you get -144?

OpenStudy (anonymous):

x=43/31 plug x value into one of the equation 5(43/31)+3y=-7 215/31+3y=-7 3y=-7-(215/31) 3y=(-217-215)/31 3y=-432/31 y=-432/(3*31) y=-144/31

OpenStudy (anonymous):

Oh i see now, thank you :D

OpenStudy (anonymous):

how did yo get 217?

OpenStudy (anonymous):

\[\frac{-7}{1}-\frac{215}{31}\] lcd=31 \[\therefore ,\frac{(-7)(31)-215}{31}\]\[\frac{217-215}{31}\]

OpenStudy (anonymous):

okay so multiplty -7x31.. lol thanks :D

OpenStudy (anonymous):

yup np :D